### v1

@@ Abstract
Uncertainty estimation is essential for the safe clinical deployment of medical image segmentation systems, enabling the identification of unreliable predictions and supporting human oversight. While prior work has largely focused on pixel-level uncertainty, landmark-based segmentation offers inherent topological guarantees yet remains underexplored from an uncertainty perspective. In this work, we study uncertainty estimation for anatomical landmark-based segmentation on chest X-rays. Inspired by hybrid neural network architectures that combine standard image convolutional encoders with graph-based generative decoders, and leveraging their variational latent space, we derive two complementary measures: (i) latent uncertainty, captured directly from the learned distribution parameters, and (ii) predictive uncertainty, obtained by generating multiple stochastic output predictions from latent samples. Through controlled corruption experiments we show that both uncertainty measures increase with perturbation severity, reflecting both global and local degradation. We demonstrate that these uncertainty signals can identify unreliable predictions by comparing with manual ground-truth, and support out-of-distribution detection on the CheXmask dataset. More importantly, we release CheXmask-U (dataset), a large scale dataset of 657,566 chest X-ray landmark segmentations with per-node uncertainty estimates, enabling researchers to account for spatial variations in segmentation quality when using these anatomical masks. Our findings establish uncertainty estimation as a promising direction to enhance robustness and safe deployment of landmark-based anatomical segmentation methods in chest X-ray. A fully working interactive demo of the method is available at CheXmask-U-demo and the source code at CheXmask-U-code.
@@
editors: Under Review for MIDL 2026
@@ 1 Introduction
In this work, we study different approaches to produce uncertainty estimates in variational architectures for landmark-based segmentation. Specifically, we focus on lung and heart segmentation in chest X-rays and exploit the VAE latent space to quantify uncertainty in two complementary ways: (i) by analyzing the latent distribution to capture model (epistemic) uncertainty, and (ii) through Monte Carlo sampling from the latent posterior, generating multiple stochastic landmark predictions per input. Recent studies have highlighted that uncertainty estimates derived from standard VAEs are not necessarily informative about true data ambiguity. In particular [catoni2025] showed that conventional VAEs may produce unreliable uncertainty representations that fail to correlate with perceptual or semantic variability. These findings motivate a closer examination of how latent-space uncertainty behaves in structured medical tasks, such as anatomical landmark localization, where meaningful confidence estimates are critical. To this end, we validate our approach through systematic corruption experiments, including occlusions and Gaussian noise, following the evaluation protocol proposed by [catoni2025], and demonstrate its utility for downstream tasks such as error prediction and out-of-distribution detection.
@@
1.1 Related work
Landmark-based segmentation. A growing body of work has explored landmark-based representations for medical structures, motivated by computational efficiency, reduced annotation requirements, and the ability to capture key anatomical relationships. Earlier shape-based approaches like statistical shape models [cootes1995active, heimann2009] encoded anatomical priors explicitly through landmark constraints but relied on hand-crafted features and limited flexibility to model large anatomical variability. Recent deep learning methods have made landmark-based segmentation more robust and scalable. Along this direction, HybridGNet [hybridgnet-miccai, hybridgnet-tmi] demonstrated the effectiveness of combining CNNs and GCNNs for decoding anatomical landmarks, effectively learning structured representations that preserve topology and spatial coherence as opposed to dense pixel-level segmentation models.
@@ 1.1 Related work
Uncertainty estimation. Uncertainty estimation in deep learning encompasses two primary types: aleatoric uncertainty (arising from inherent data noise and annotation ambiguity) and epistemic uncertainty (stemming from model limitations and insufficient training data [kendall2017]). In medical applications, both sources contribute to prediction uncertainty and must be carefully accounted for to ensure reliable and trustworthy clinical decision support. This motivates the need for robust uncertainty quantification (UQ) frameworks that can explicitly assess model confidence and guide interpretation.
@@ 2.2 Uncertainty Estimation
    Latent distribution analysis (epistemic uncertainty). Given an encoded latent vector for a single input image, its variance $\sigma^{2}$ captures model (epistemic) uncertainty regarding landmark placement, providing a global measure of confidence for the predicted anatomical configuration.
@@ 3.1 Dataset, implementation and training details
We trained a HybridGNet model for lung and heart landmark-based segmentation using the chest X-ray datasets JSRT [shiraishi2000development], Padchest [bustos2020padchest], Montgomery [montgomeryset] and Shenzhen [shenzhen]. Graphs were constructed from the landmark coordinates, with each landmark defining a node and the adjacency matrix encoding anatomical connectivity, shared across all subjects. Following the heterogeneous-label training strategy of [hybridgnet-isbi], each batch contains images from a single dataset, and the loss is evaluated only on the nodes corresponding to available annotations. Compared to the original model, we used a higher weighting for the KL divergence term to encourage more structured latent representations: the KL weight was initialized at $1\times 10^{-5}$ and gradually increased to $1\times 10^{-2}$ during training. On an NVIDIA TITAN Xp GPU, encoding takes 15ms per image for both variants. Decoding is efficient at 7.4-19.7ms (faster without skip connections), as multiple latent samples can be batched together. This makes encoding the main bottleneck, as it is performed only once per image. We now describe the different experiments that validate our uncertainty quantification approach.
@@
Validating UQ under Image Occlusions. We first simulated occlusions under the hypothesis that uncertainty should be higher in occluded areas, applying artificial black-square masks on different image regions. For each image, node-level predictive uncertainty was computed from $N=50$ stochastic output predictions $\{\hat{X}^{(i)}\}_{i=1}^{N}$ using the model with skip-connections. As shown in Figure 2, nodes located in the occluded regions exhibit high uncertainty compared to unaffected regions, reflecting the model’s ability to localize confidence degradation under partial information loss. Moremover, the boxplot shows uncertainty for nodes falling under occluded and visible areas, confirming significantly larger values for the occluded ones.
@@ 3.2 The CheXmask-U Dataset
Validation. We validated the per-node predictive uncertainty against actual landmark error, derived from 255 manually annotated images, with two independent annotations. Figure 4 shows a strong positive correlation between predicted uncertainty and actual error, confirming that higher uncertainty reliably indicates less accurate predictions. We further validated CheXmask-U uncertainty estimates by evaluating their relationship with RCA-estimated Dice scores. For each image, the mean node-wise standard deviation was computed as a proxy for overall uncertainty. Figure 3(c) shows a clear anti-correlation: images with higher predicted uncertainty tend to have lower RCA-DSC. Importantly, CheXmask-U provides node-level granularity that RCA-Dice cannot: an image may have acceptable overall quality but contain specific anatomical regions with high uncertainty—information that is crucial for region-specific downstream applications. By providing both landmark coordinates and per-node uncertainty estimates, CheXmask-U enables researchers to: (i) selectively use landmarks based on confidence thresholds for their specific anatomical regions of interest, (ii) weight contributions from different anatomical regions based on reliability, and (iii) assess segmentation quality at a spatial granularity not available from image-level scores alone. All uncertainty estimates are pre-computed and publicly available, eliminating the need for users to perform uncertainty quantification themselves.
@@ 4 Conclusion
Our experiments demonstrate that both latent-space and node-wise predictive uncertainties respond appropriately to image perturbations, including occlusions and Gaussian noise, and serve as effective indicators of out-of-distribution chest X-rays. These validations confirm that landmark-based uncertainty captures meaningful information about prediction reliability.

### v2

@@ Abstract
In this work, we study uncertainty estimation for anatomical landmark-based segmentation on chest X-rays. Inspired by hybrid neural network architectures that combine standard image convolutional encoders with graph-based generative decoders, and leveraging their variational latent space, we derive two complementary measures: (i) latent uncertainty, captured directly from the learned distribution parameters, and (ii) predictive uncertainty, obtained by generating multiple stochastic output predictions from latent samples. Through controlled corruption experiments we show that both uncertainty measures increase with perturbation severity, reflecting both global and local degradation. We demonstrate that these uncertainty signals can identify unreliable predictions by comparing with manual ground-truth, and support out-of-distribution detection on the CheXmask dataset. More importantly, we release CheXmask-U (dataset), a large scale dataset of 657,566 chest X-ray landmark segmentations with per-node uncertainty estimates, enabling researchers to account for spatial variations in segmentation quality when using these anatomical masks. Our findings establish uncertainty estimation as a promising direction to enhance robustness and safe deployment of landmark-based anatomical segmentation methods in chest X-ray. A fully working interactive demo of the method is available at CheXmask-U-demo and the source code at CheXmask-U-code.
@@
editors: Accepted for publication at MIDL 2026
@@ 1 Introduction
In this work, we study different approaches to produce uncertainty estimates in variational architectures for landmark-based segmentation. Specifically, we focus on lung and heart segmentation in chest X-rays and exploit the VAE latent space to quantify uncertainty in two complementary ways: (i) by analyzing the latent distribution to capture latent-variable–based predictive uncertainty uncertainty, and (ii) through Monte Carlo sampling from the latent posterior, generating multiple stochastic landmark predictions per input. Recent studies have highlighted that uncertainty estimates derived from standard VAEs are not necessarily informative about true data ambiguity. In particular [catoni2025] showed that conventional VAEs may produce unreliable uncertainty representations that fail to correlate with perceptual or semantic variability. These findings motivate a closer examination of how latent-space uncertainty behaves in structured medical tasks, such as anatomical landmark localization, where meaningful confidence estimates are critical. To this end, we validate our approach through systematic corruption experiments, including occlusions and Gaussian noise, following the evaluation protocol proposed by [catoni2025], and demonstrate its utility for downstream tasks such as error prediction and out-of-distribution detection.
@@
2 Related work
2.1 Landmark-based segmentation.
A growing body of work has explored landmark-based representations for medical structures, motivated by computational efficiency, reduced annotation requirements, and the ability to capture key anatomical relationships. Earlier shape-based approaches like statistical shape models [cootes1995active, heimann2009] encoded anatomical priors explicitly through landmark constraints but relied on hand-crafted features and limited flexibility to model large anatomical variability. Recent deep learning methods have made landmark-based segmentation more robust and scalable. Along this direction, HybridGNet [hybridgnet-miccai, hybridgnet-tmi] demonstrated the effectiveness of combining CNNs and GCNNs for decoding anatomical landmarks, effectively learning structured representations that preserve topology and spatial coherence as opposed to dense pixel-level segmentation models.
@@ 2 Related work
2.2 Uncertainty estimation.
Uncertainty estimation in deep learning encompasses two primary types: aleatoric uncertainty (arising from inherent data noise and annotation ambiguity) and epistemic uncertainty (stemming from model limitations and insufficient training data [kendall2017]). In medical applications, both sources contribute to prediction uncertainty and must be carefully accounted for to ensure reliable and trustworthy clinical decision support. This motivates the need for robust uncertainty quantification (UQ) frameworks that can explicitly assess model confidence and guide interpretation. In this work, we primarily focus on uncertainty derived from latent-variable sampling in a variational model, without marginalizing over model parameters.
@@ 3.2 Uncertainty Estimation
    Latent distribution analysis (latent-variable uncertainty). Given an encoded latent vector for a single input image, its variance $\sigma^{2}$ captures model (epistemic) uncertainty regarding landmark placement, providing a global measure of confidence for the predicted anatomical configuration.
@@ 4.1 Dataset, implementation and training details
We trained a HybridGNet model for lung and heart landmark-based segmentation using the chest X-ray datasets JSRT [shiraishi2000development], Padchest [bustos2020padchest], Montgomery [montgomeryset] and Shenzhen [shenzhen]. Graphs were constructed from the landmark coordinates, with each landmark defining a node and the adjacency matrix encoding anatomical connectivity, shared across all subjects. Following the heterogeneous-label training strategy of [hybridgnet-isbi], each batch contains images from a single dataset, and the loss is evaluated only on the nodes corresponding to available annotations. Compared to the original model, we used a higher weighting for the KL divergence term to encourage more structured latent representations: the KL weight was initialized at $1\times 10^{-5}$ and gradually increased to $1\times 10^{-2}$ during training. We observed that very low KL weights lead to near-zero latent posterior variance, resulting in uninformative uncertainty estimates, while excessively large values degrade segmentation accuracy. The adopted KL schedule prevents the decoder from ignoring the latent variables and preserves a structured latent space with stable uncertainty estimates while maintaining predictive performance. On an NVIDIA TITAN Xp GPU, encoding takes 15ms per image for both variants. Decoding is efficient at 7.4-19.7ms (faster without skip connections), as multiple latent samples can be batched together. This makes encoding the main bottleneck, as it is performed only once per image. All experiments are conducted using $N=50$ stochastic samples, which we found to provide stable uncertainty estimates with a favorable runtime trade-off; sensitivity to $N$ and runtime analysis are reported in Appendix A. We now describe the different experiments that validate our uncertainty quantification approach.
@@
Validating UQ under Image Occlusions. We first simulated occlusions under the hypothesis that uncertainty should be higher in occluded areas, applying artificial black-square masks on different image regions. For each image, node-level predictive uncertainty was computed from $N=50$ stochastic output predictions $\{\hat{X}^{(i)}\}_{i=1}^{N}$ using the model with skip-connections. As shown in Figure 2, nodes located in the occluded regions exhibit high uncertainty compared to unaffected regions, reflecting the model’s ability to localize confidence degradation under partial information loss. Moreover, the boxplot shows uncertainty for nodes falling under occluded and visible areas, confirming significantly larger values for the occluded ones.
@@ 4.2 The CheXmask-U Dataset
Figure 5: Reliability analysis of node-level uncertainty. Landmark predictions are binned into uncertainty quantiles, and for each bin the mean observed landmark error is plotted against the mean predicted uncertainty with results shown separately for the two annotators.
@@
Validation. We validated the per-node predictive uncertainty against actual landmark error, derived from 255 manually annotated images, with two independent annotations. Figure 4 shows a strong positive correlation between predicted uncertainty and actual error, confirming that higher uncertainty reliably indicates less accurate predictions. To further assess the consistency between predicted uncertainty and observed error, we performed a reliability analysis by binning landmark predictions according to their predicted uncertainty. Specifically, nodes were grouped into uncertainty quantiles, and for each bin we computed the mean predicted uncertainty and the corresponding mean landmark error. Figure 5 shows that bins with higher predicted uncertainty exhibit systematically larger observed errors for both annotators.
We further validated CheXmask-U uncertainty estimates by evaluating their relationship with RCA-estimated Dice scores. For each image, the mean node-wise standard deviation was computed as a proxy for overall uncertainty. Figure 3(c) shows a clear anti-correlation: images with higher predicted uncertainty tend to have lower RCA-DSC. Importantly, CheXmask-U provides node-level granularity that RCA-Dice cannot: an image may have acceptable overall quality but contain specific anatomical regions with high uncertainty—information that is crucial for region-specific downstream applications. By providing both landmark coordinates and per-node uncertainty estimates, CheXmask-U enables researchers to: (i) selectively use landmarks based on confidence thresholds for their specific anatomical regions of interest, (ii) weight contributions from different anatomical regions based on reliability, and (iii) assess segmentation quality at a spatial granularity not available from image-level scores alone. All uncertainty estimates are pre-computed and publicly available, eliminating the need for users to perform uncertainty quantification themselves.
Figure 6: Correlation between predictive uncertainty and segmentation error (1-Dice) for different uncertainty quantification methods at the pixel level. Ground-truth masks are obtained by averaging the annotations from two experts.
4.3 Comparison with pixel-level UQ baselines
We compare our landmark-based method against established pixel-level uncertainty quantification methods under a unified evaluation protocol. We consider the following approaches: (i) a pixel-level U-Net segmentation model with Monte Carlo (MC) Dropout, (ii) the probabilistic segmentation model PHiSeg [phiseg], (iii) HybridGNet with MC Dropout, and (iv) the proposed variational HybridGNet. For all methods, uncertainty is computed in the same manner: we generate $N=50$ stochastic pixel-level segmentation masks per image, compute the pixel-wise predictive entropy of the mean segmentation, and average entropy over ground-truth foreground pixels, following [nair2019exploring]. This yields a single global uncertainty score per image that can be directly compared across methods.
To implement HybridGNet with MC Dropout, we remove the variational latent space and train a deterministic version of HybridGNet with dropout applied to the encoder layers. At inference time, dropout is kept active to produce multiple stochastic predictions. For both HybridGNet variants, landmark-based outputs are deterministically converted into pixel-level segmentation masks (by filling-in the contours) before computing entropy, ensuring a fair comparison with pixel-based methods. Figure 6 reports the correlation between predictive entropy and segmentation error (1 - Dice) across methods. The proposed variational HybridGNet exhibits a stronger uncertainty–error correlation than the pixel-level baselines, indicating a more reliable global uncertainty signal. This analysis complements the node-level uncertainty evaluation presented in the main paper, which provides fine-grained spatial information that cannot be obtained from pixel-based UQ methods alone.
Finally, we note a practical advantage of the proposed approach. Unlike MC Dropout or pixel-level variational models, which require a full forward pass for each stochastic sample, our method encodes the image only once and performs multiple stochastic decodings from the latent space, resulting in significantly lower computational overhead for uncertainty estimation.
@@ 5 Conclusion
Our experiments demonstrate that both latent-space and node-wise predictive uncertainties respond appropriately to image perturbations, including occlusions and Gaussian noise, and serve as effective indicators of out-of-distribution chest X-rays. These validations confirm that landmark-based uncertainty captures meaningful information about prediction reliability. In addition, comparisons with established pixel-level uncertainty methods show that the proposed approach provides complementary, spatially informative signals with lower computational overhead.
@@
[…]
EF was supported by the Google Award for Inclusion Research, a Googler Initiated Grant and the Distinguished International Associate award of the Royal Academy of Engineering.
Appendix A Sensitivity to number of samples
We analyze the sensitivity of the proposed uncertainty estimates to the number of samples used at inference time. We report the correlation between predicted uncertainty and landmark error, measured against inter-annotator discrepancies from two independent expert annotations, for different values of $N$. Table 1 shows that uncertainty–error agreement remains stable across a wide range of sample counts, with performance saturating well before $N=50$. This indicates that the uncertainty estimates are robust to the choice of $N$ and that a moderate number of samples is sufficient in practice.
Model NN Annotator 1 Annotator 2 Runtime (s)
HybridGNet (MC Dropout) 10 0.52 0.52 32.2
25 0.54 0.54 74.0
50 0.59 0.59 142.3
100 0.59 0.60 281.4
200 0.59 0.59 556.8
HybridGNet (Variational) 10 0.53 0.55 7.2
25 0.56 0.57 8.0
50 0.59 0.60 9.0
100 0.60 0.60 11.5
200 0.60 0.61 16.4
Table 1: Sensitivity of uncertainty–error correlation to the number of Monte Carlo samples NN for variational HybridGNet and HybridGNet with MC Dropout. Correlation is measured between predicted uncertainty and landmark error using inter-annotator discrepancies from two independent expert annotations. Total runtime over all 255 annotations is reported in each case.